# A Comparative Study on State-Action Spaces for Learning Viewpoint Selection and Manipulation with Diffusion Policy

### v2

@@ IV-B Look-At IK
Building on the dual-arm IK solver, the look-at IK solver is introduced to further enhance viewpoint selection. It incorporates an additional viewpoint-selection loss, which automatically calculates the orientation of the viewpoint end-effector. This reduces the state-action space to a look-at space variant $\mathcal{L}_{\text{E}},\mathcal{L}_{\text{Q}},\mathcal{L}_{\text{AA}}$, which has lower dimensionality compared to the full end-effector space variants $\mathcal{E}_{\text{E}},\mathcal{E}_{\text{Q}},\mathcal{E}_{\text{AA}}$.
@@
The viewpoint orientation objective term measures the distance between the visual target $\mathbf{p_{t}}$ and its projection onto a line segment defined by the viewpoint position $\mathbf{p_{v}}$ and a distant point $\mathbf{p_{f}}$, located along the $z$ axis direction $\mathbf{w}_{z}$ at a fixed distance $\delta$:
@@
In this framework, the visual target $\mathbf{p_{t}}$ is set as a fixed offset relative to the manipulation arm’s end-effector position $\mathbf{p_{m}}$. Finally, the entire objective function for the look-at IK solver becomes:
@@
IV-C Diffusion Policy
In this work, we use the diffusion policy [5] as the visuomotor policy for imitation learning. We use U-Net [30] as the backbone for the diffusion policy. The diffusion policy employs receding horizon control and models the conditional distribution $p(\mathbf{A}_{t}\mid\mathbf{O}_{t})$. During inference, the diffusion policy starts from Gaussian noise $\mathbf{A}_{t}^{K}$, where $K$ is the number of steps for denoising. The denoising process repeatedly utilizes the following equation:
until the process reaches the output $\mathbf{A}_{t}^{0}$ of the policy.
@@ V-A Implementation
We compared diffusion policies trained with different state-action spaces in various simulated and real-world tasks with demonstrations collected from teleoperation and the same hyperparameters as [5]. For the look-at IK, we set $\delta$ to 999 and adjusted $p_{t}$ to $p_{t}-[0.0,0.0,-0.15]^{T}$ with $m$ as the unit. We used PyTorch [26] as the machine learning framework for training the policy. All IK optimizations are performed using SLSQP [18] in NLopt [15]. Both training and inference were performed using an AMD PRO 5975WX CPU, dual 4090 GPUs, and 128GB RAM. The primary metrics throughout all evaluations were task success rate and the mean and standard deviation of task completion duration.
@@
Simulated Pick and Place (200 Epochs) $\mathcal{C}$ N/A N/A 0.00
@@
$\mathcal{L}_{\text{AA}}$ 498.55 stp 13.47 stp 0.05
Simulated Microwave (10 Epochs) $\mathcal{C}$ 446.98 stp 93.83 stp 0.30
@@
Real-World Block Stacking (300 Epochs) $\mathcal{C}$ N/A N/A 0.00
@@
Real-World Drawer Interaction (300 Epochs) $\mathcal{C}$ N/A N/A 0.00
@@ V-B Simulation Experiments
The simulation environments, developed using MuJoCo [38], include a pick-and-place scenario and a microwave scenario. In the pick-and-place scenario, the agent must pick up a green cube and place it on a red goal position, with the cube randomly generated within a 15 cm radius around the goal. This scenario is shown in the first row of Fig. 3. The microwave scenario, depicted in the second row of Fig. 3, requires the agent to open the microwave door by at least 0.5 radians. Results for all simulation experiments were calculated from 100 inference rollouts. The pick-and-place policy was trained with 282 rollouts, and the microwave policy with 151 rollouts.
The first and second rows in Tab. I present the results for the pick-and-place and microwave scenarios, respectively, measured in steps (stp) for task duration. Overall, state-action spaces using Euler angles performed better compared to other rotation representations. Using look-at IK improved success rates for all tasks, likely due to reduced dimensionality. Policies with look-at IK also had lower mean and standard deviation in task duration, indicating less uncertainty. However, differences in dimensionality cannot explain why $\mathcal{L}_{\text{AA}}$ that has less amount of elements compared to $\mathcal{E}_{\text{E}}$ and $\mathcal{E}_{\text{Q}}$ performs worse in the pick and place scenario.
The microwave task was easier to learn than the pick-and-place task. For example, the policy trained with $\mathcal{L}_{\text{E}}$ reached a success rate of 1.00 after 10 training epochs, while the pick-and-place task reached 0.92 after 200 epochs despite having a larger demonstration dataset. Given the ease of the microwave task, we explored whether additional training epochs could help policies from different state-action spaces achieve comparable performance. As shown in Fig. 4, increasing training epochs led to success rates of policies from different state-action spaces becoming more comparable, indicating that performance differences are likely due to the ability of the network to capture patterns in the corresponding space.
@@ V-C Real-World Experiments
We conducted real-world experiments in two scenarios: block stacking using 3D-printed objects from [19] and drawer interaction with a 3D-printed drawer from [13]. The block stacking task involved non-cuboid objects, requiring precise gripper orientation for successful grasping, while the drawer interaction task required the agent to get rid of visual occlusion from a cardboard box, place a red target cube into the drawer, and close it. The block stacking policy was trained with 201 rollouts, and the drawer interaction policy with 67 rollouts. Each task was evaluated with 30 inference rollouts. As shown in the third and fourth rows of Tab. I, the use of the look-at IK solver further improved both success rate and performance stability, and state-action spaces using Euler angles or quaternions for rotation representation generally achieved higher success rates, further indicating that dimensionality is not the primary factor for affecting learning performance given that quaternions and axis-angle representations have the same amount of elements.
@@ V-D Frequency Analysis
Figure 5: Frequency heatmaps of concatenated state and action arrays from demonstration datasets across different evaluation scenarios (PnP for pick-and-place, M for microwave, BS for block stacking, DI for drawer interaction). The horizontal axes indicate the lowest and highest frequency indices, while the vertical axes correspond to different state and action spaces. Color intensity denotes the logarithmically scaled magnitude of the frequency components.
To further investigate the performance differences among state-action spaces, we conducted a frequency analysis inspired by recent work on the interplay between frequency and CNNs [43]. We performed a Fast Fourier Transform to convert the time-domain signals into the frequency domain and calculated the logarithm of the magnitudes of the Fourier coefficients:
@@
The results, shown in Fig. 5, reveal a stronger correlation between success rates and the magnitude of HFC compared to dimensionality. For example, the relative ease of learning with $\mathcal{L}_{\text{E}}$ may be due to the low magnitude of HFC, whereas the difficulty when learning with $\mathcal{E}_{\text{AA}}$ may be attributed to the high magnitude of HFC despite having the same dimensionality as $\mathcal{E}_{\text{Q}}$.
@@ VI Discussion
VI-A Implications
Learning from state-action spaces with Euler angles as the rotation representation yielded the best performance in most of our evaluation tasks. However, Euler angles are known to have several disadvantages in both spatial representation and optimization, such as gimbal lock due to singularities [35] and a non-smooth optimization landscape due to discontinuities [11]. Despite these drawbacks, the HFC present in other rotation representations and configuration spaces pose challenges for current diffusion policy backbones, especially when dealing with high-dimensional state-action spaces and long task horizons.
Previous work has shown that while CNNs are effective at capturing HFC, the use of residual structures can suppress HFC learning [43]. Removing these residuals from the UNet architecture, which is commonly used in diffusion policies, is not a viable solution as they are critical for deep network training as a mitigation for vanishing gradients [12]. Although Transformers [44] are another popular choice for generative diffusion models [27], they are better suited for capturing LFC rather than HFC [25], making them less ideal for diffusion policies with state-action spaces that contain significant HFC.
VI-B Limitations and Future Work
Despite the growing interest in the impact of different state-action spaces and rotation representations on learning performance and data efficiency, this direction remains underexplored, especially concerning the potential connection between frequency and model architecture. Although our work proposes the hypothesis and provides empirical evidence that HFC affects learning performance for current diffusion policies, more thorough theoretical analysis is needed to further explore the interplay between frequency and model architecture.
Given the lack of suitable network architectures for learning from state-action spaces with significant HFC, and considering the optimization and spatial representation benefits of these spaces, another future direction is to propose new diffusion backbones capable of capturing HFC from these spaces, potentially improving sample efficiency and generalizability. Incorporating more geometrical inductive biases into the design of network architectures may be necessary. Besides improving on architecture, another future direction would be projecting datasets with intense HFC to higher dimensional space, similar to how [23] dealing with HFC for better capturing details in learning radiance field. Additionally, this work sets the visual target $\mathbf{p_{t}}$ as a fixed offset from the end-effector position of the manipulation arm $\mathbf{p_{m}}$. Future work will also explore leveraging semantic information from the environment to set $\mathbf{p_{t}}$ based on detected semantic objects in the environment.

### v3

@@ IV-B Look-At IK
Building on the traditional end-effector IK solver, the look-at IK solver is introduced to further enhance viewpoint selection. It incorporates an additional viewpoint-selection loss, which automatically calculates the orientation of the viewpoint end-effector. This reduces the state-action space to a look-at space variant $\mathcal{L}_{\text{E}},\mathcal{L}_{\text{Q}},\mathcal{L}_{\text{AA}}$, which has lower dimensionality compared to the full end-effector space variants $\mathcal{E}_{\text{E}},\mathcal{E}_{\text{Q}},\mathcal{E}_{\text{AA}}$.
@@
The viewpoint orientation objective term measures the distance between the visual target $\mathbf{p_{t}}$ and its projection onto a line segment defined by the viewpoint position $\mathbf{p_{v}}$ and a distant point $\mathbf{p_{f}}$, located along the $z$ axis direction $\mathbf{w}_{z}$:
@@
In practice, we set the distance factor $\delta$ to a large constant (e.g., 999) to ensure $\mathbf{p_{f}}$ is always sufficiently distant.
@@
In this framework, the visual target $\mathbf{p_{t}}$ is set as the gripper position $\mathbf{p_{m}}$. Finally, the entire objective function for the look-at IK solver becomes:
@@ V-A Implementation
To demonstrate the effectiveness of our proposed method, we compared diffusion policies trained within different state-action spaces in various simulated and real-world tasks with demonstrations collected from teleoperation and the same hyperparameters as [5]. These baselines include the End-Effector Quaternion space ($\mathcal{E}_{\text{Q}}$), which aligns with the approach used in concurrent work on active perception and manipulation [43], allowing for a direct comparison.
We used PyTorch [22] as the machine learning framework for training the policy. All IK optimizations are performed using SLSQP [16] in NLopt [13]. Both training and inference were performed using an AMD PRO 5975WX CPU, dual 4090 GPUs, and 128GB RAM. The primary metrics throughout all evaluations were task success rate and the mean and standard deviation of task completion duration.
@@
Simulated
Pick and Place
(200 Epochs)
$\mathcal{C}$ N/A N/A 0.00
@@
$\mathcal{L}_{\text{AA}}$ 498.55 stp 13.47 stp 0.21
Simulated
Microwave
(10 Epochs)
$\mathcal{C}$ 446.98 stp 93.83 stp 0.30
@@
Real-World
Block Stacking
(300 Epochs)
$\mathcal{C}$ N/A N/A 0.00
@@
Real-World
Drawer Interaction
(300 Epochs)
$\mathcal{C}$ N/A N/A 0.00
@@ V-B Simulation Experiments
The simulation environments for dual-arm active perception and manipulation, developed using MuJoCo [35], include a pick-and-place scenario and a microwave scenario. In the pick-and-place scenario, the agent must pick up a green cube and place it on a red goal position, with the cube randomly generated within a 15 cm radius around the goal. This scenario is shown in the first row of Fig. 3. The microwave scenario, depicted in the second row of Fig. 3, requires the agent to open the microwave door by at least 0.5 radians. Results for all simulation experiments were calculated from 100 inference rollouts. The pick-and-place policy was trained with 282 rollouts, and the microwave policy with 151 rollouts.
The simulation results in Tab. I (row 1-2) show that our framework, which leverages the look-at IK, consistently and significantly outperforms the baselines. Diffusion policies trained in the look-at spaces achieve higher success rates—up to 4.5 times greater than their end-effector counterparts—and exhibit more stable performance with lower variance in task duration.
This performance gain, however, cannot be attributed solely to dimensionality reduction. We observe two counter-intuitive results: first, there are instances where policies in lower-dimensional spaces underperform those in higher-dimensional ones—for example, the policy in $\mathcal{L}_{\text{AA}}$ (11-dim) underperforms the one in $\mathcal{E}_{\text{E}}$ (13-dim)—contradicting the common assumption that reducing dimensionality simplifies the learning problem and improves performance; second, representations using Euler angles consistently outperform theoretically superior quaternions and axis-angles, despite their known issues with singularities and discontinuities [32, 10]. These findings strongly suggest that a different, underlying factor is at play, motivating our frequency analysis in §V-D.
The microwave task was easier to learn than the pick-and-place task. For example, the policy trained with $\mathcal{L}_{\text{E}}$ reached a success rate of 1.00 after 10 training epochs, while the pick-and-place task reached 0.92 after 200 epochs despite having a larger demonstration dataset. Given the ease of the microwave task, we explored whether additional training epochs could help policies from C-Space and with end-effector IK achieve comparable performance. As shown in Fig. 4, increasing training epochs led to success rates of baseline policies becoming more comparable, indicating that performance differences are likely due to better training efficiency of our proposed framework. This also indicates that diffusion policies can capture the data pattern easier with look-at IK compared to baselines.
@@ V-C Real-World Experiments
We conducted real-world experiments in two scenarios: block stacking using 3D-printed objects from [17] and drawer interaction with a 3D-printed drawer from [11]. The block stacking task involved non-cuboid objects, requiring precise gripper orientation for successful grasping, while the drawer interaction task required the agent to get rid of visual occlusion from a cardboard box, place a red target cube into the drawer, and close it. The block stacking policy was trained with 201 rollouts, and the drawer interaction policy with 67 rollouts. Each task was evaluated with 30 inference rollouts. As shown in Tab. I (rows 3-4), policies learned with our look-at IK again yields substantial improvements in success rate and performance stability across both real-world tasks. The results also reinforce that dimensionality is not the primary performance driver, as policies trained in quaternion and axis-angle spaces—which have the same dimensionality—show markedly different success rates.
@@ V-D Frequency Analysis
Figure 5: Frequency heatmaps of concatenated state and action arrays from demonstration datasets across different evaluation scenarios (PnP for pick-and-place, M for microwave, BS for block stacking, DI for drawer interaction). Min and Max ticks indicate the lowest and highest frequency components. Color intensity denotes the logarithmically scaled magnitude of the frequency components.
To further investigate the performance improvement when training diffusion policies with our proposed look-at IK, we conducted a frequency analysis inspired by recent work on the interplay between frequency and CNNs [40]. We performed a Fast Fourier Transform to convert the time-domain signals into the frequency domain and calculated the logarithm of the magnitudes of the Fourier coefficients:
@@
The results, shown in Fig. 5, reveal that policy success rate is more strongly correlated with the magnitude of HFC than with space dimensionality. Notably, our proposed look-at spaces consistently exhibit lower HFC magnitudes across all tasks when compared to the baseline C-Space and end-effector spaces. This analysis also explains the surprising effectiveness of Euler angles, as they consistently produce lower HFC than other rotation representations, while axis-angle representations exhibit the highest. This finding aligns with machine learning literature showing that common diffusion backbones, which rely on residual connections, are less effective at learning from HFC [40, 26]. Therefore, by creating a state-action space with naturally lower HFC, our framework makes the learning problem more tractable for the diffusion policy, leading to the superior performance and efficiency observed in our experiments.
@@ VI Discussion
In this work, we introduced a framework that integrates a diffusion policy with a novel Look-At IK solver to address the challenges of manipulation with active perception. While concurrent work has approached this problem with standard end-effector IK [43], our method infuses a structural prior into the system. By automating viewpoint orientation, our framework creates a simplified space that is easier to learn for diffusion policies.
Beyond the immediate performance gains, our analysis provides a deeper insight by linking this success to the frequency characteristics of the space. We provide empirical evidence that the difficulty faced by standard approaches is strongly correlated with HFC, a factor not yet widely considered in the robotics community for this type of task. We hope this analysis encourages the community to consider robot learning problems through a new lens.
While our Look-At IK solver effectively mitigates the HFC issue for active perception, it is a task-specific solution. The fundamental challenge—that common diffusion backbones struggle to learn HFC—remains a general problem for many robotics tasks. Therefore, a critical direction for future work is to design new diffusion backbones that can capture HFC more effectively.